### v1

# Planetarium[Uncaptioned image]: A Rigorous Benchmark for Translating Text to Structured Planning Languages

Abstract
Many recent works have explored using language models for planning problems. One line of research focuses on translating natural language descriptions of planning tasks into structured planning languages, such as the planning domain definition language (PDDL). While this approach is promising, accurately measuring the quality of generated PDDL code continues to pose significant challenges. First, generated PDDL code is typically evaluated using planning validators that check whether the problem can be solved with a planner. This method is insufficient because a language model might generate valid PDDL code that does not align with the natural language description of the task. Second, existing evaluation sets often have natural language descriptions of the planning task that closely resemble the ground truth PDDL, reducing the challenge of the task. To bridge this gap, we introduce Planetarium, a benchmark designed to evaluate language models’ ability to generate PDDL code from natural language descriptions of planning tasks. We begin by creating a PDDL equivalence algorithm that rigorously evaluates the correctness of PDDL code generated by language models by flexibly comparing it against a ground truth PDDL. Then, we present a dataset of $132,037$ text-to-PDDL pairs across $13$ different tasks, with varying levels of difficulty. Finally, we evaluate several API-access and open-weight language models that reveal this task’s complexity. For example, $87.6\%$ of the PDDL problem descriptions generated by GPT-4o are syntactically parseable, $82.2\%$ are valid, solve-able problems, but only $35.1\%$ are semantically correct, highlighting the need for a more rigorous benchmark for this problem.
1 Introduction
Recently, there has been growing interest in using large language models (LLMs) to solve planning problems. Some research has focused on generating plans directly with LLMs Valmeekam et al. (2023a, 2022, b); Silver et al. (2022, 2023). However, this approach has shown limited success; GPT-4 only achieves $35\%$ accuracy on simple planning problems Valmeekam et al. (2023a). Another line of research uses LLMs to convert natural language prompts into structured planning languages, such as the planning domain definition language (PDDL) Liu et al. (2023); Xie et al. (2023b); Guan et al. (2023); Chalvatzaki et al. (2023); Yang et al. (2023b). Early evidence suggests this method performs better than generating plans directly with LLMs Liu et al. (2023). Despite its promise, there is a lack of rigorous techniques and benchmarks for evaluating the translation of natural language planning descriptions to PDDL.
Translating natural language to PDDL enables a hybrid, best-of-both-worlds approach. The LLM is responsible for interpreting natural language, and the resulting PDDL can be given to traditional, symbolic planners that have been developed over decades Fikes and Nilsson (1971); McDermott et al. (1998); Helmert (2006); Baier et al. (2009); Vidal and Geffner (2006). These traditional planners are efficient and ensure the correctness of their solutions. In contrast, leaving the planning to LLMs does not guarantee correctness, making them unreliable for critical applications. Despite this shortcoming, LLMs are an enticing approach, as using traditional planners requires domain expertise and expertise in modeling planning problems. By using LLMs to translate natural language into PDDL, we can better leverage the strengths of existing planners.
Using LLMs to translate natural language to PDDL is an instance of a code generation task. Evaluating code generation tasks in general is highly challenging. Some benchmarks for code generation use match-based metrics that look for segments that overlap with ground truth code Papineni et al. (2002); Lin (2004); Ren et al. (2020), but match-based metrics cannot cover the vast space of equivalent programs Chen et al. (2021). Therefore, benchmarks often test functional correctness using a suite of unit tests Roziere et al. (2020); Kulal et al. (2019). For planning problems, existing works use “plan validators” to check if the generated code can be solved with a traditional planner Liu et al. (2023); Silver et al. (2022, 2023); Guan et al. (2023). We argue that these validators alone are insufficient to determine if PDDL generation is correct. This is because the LLM can generate valid PDDL that has nothing to do with the user’s instructions and still be considered correct.
Figure 1: An illustration of how multiple PDDL problem descriptions can describe the same planning problem. All four PDDL problem descriptions represent the planning problem of stacking one block onto another.
A rigorous evaluation of LLMs as generators of structured planning languages requires a precise definition of what it means for generated code to be correct, and an efficient, automatic way of checking it. This requirement is hard because many instances of PDDL can represent the same planning problem, but it is not always obvious when they are equivalent (Figure 1). Properly checking PDDL equivalence requires symbolic interpretation.
To address this challenge, we introduce Planetarium, a benchmark to evaluate LLMs on translating natural language descriptions of planning problems into PDDL. Our contributions are as follows:
  • Rigorous Evaluation of PDDL Equivalence. We formally define planning problem equivalence and create an algorithm for checking whether two PDDL problems satisfy this definition. This algorithm transforms the PDDL code into scene graphs, computes an expansion of the goal states for both PDDL problems, and then performs isomorphism checks between the graphs. Our method ensures two PDDL problems match if and only if they represent the same underlying planning problem. Further, we show how to make this algorithm practically efficient for two domains: Blocks World and Gripper McDermott (2000) (Appendix C).
  • Benchmark Data for PDDL Generation. We present a dataset based on the International Planning Competition (IPC) McDermott (2000); Vallati et al. (2015); Taitler et al. (2024); Seipp et al. (2022), crafting $132,037$ ground truth PDDL problems and corresponding text descriptions capturing a range of planning problems. Each task varies in two dimensions to assess the difficulty of the PDDL generation task: abstraction and size.
  • Broad Evaluation of Current LLMs. Finally, we evaluate a range of API-access and open-access LLMs on Planetarium. We evaluate in both a zero-shot setting and after fine-tuning. We find that this task is very challenging. GPT-4o in a zero-shot setting gets only $35.1\%$ correct. Instances with abstract descriptions or many propositions are particularly challenging. Planetarium can therefore serve as a benchmark of progress on this important problem. To enable future development and evaluation of LLMs we release all the code and data for Planetarium at https://github.com/BatsResearch/planetarium.
2 Related Work
LLMs as Planners.
There is growing interest in using LLMs for solving planning problems. Several works have demonstrated the potential of using LLMs for decision making processes Sharma et al. (2022); Ren et al. (2023); Ichter et al. (2022); Singh et al. (2022). Some techniques that enable the use of LLMs as planners involve decomposing abstract tasks into concrete, atomic executable actions for an agent to perform Huang et al. (2022b, a); Sharma et al. (2022). Other approaches generate plan actions by scoring the possible next steps in a sequence of actions Ichter et al. (2022); Ren et al. (2023). When using LLMs to generate plans directly, the LLM is given a natural language representation of a planning problem with the goal of generating a plan. We refer to this line of work as “LLMs as planners.” One of the main findings is that LLMs have limited ability to generate and validate plans on their own, even for simple planning tasks Valmeekam et al. (2022); Silver et al. (2022); Valmeekam et al. (2023b).
Planner-Augmented LLMs. Given LLMs’ poor performance in classical planning tasks Valmeekam et al. (2023a, 2022); Silver et al. (2022), new approaches extend LLMs by incorporating classical planning techniques Liu et al. (2023); Guan et al. (2023). Some techniques frame the problem as a machine translation task, transforming a natural language text description of a planning problem into PDDL to use with classical planners Liu et al. (2023); Guan et al. (2023); Xie et al. (2023a). We refer to this line of work as “Planner-Augmented LLMs.” Other similar approaches translate natural language into alternative representations of the planning problem, such as finite state automata Yang et al. (2023a) or logic programming Yang et al. (2023b), to solve them.
Benchmarking LLMs on Planning Tasks. Significant efforts have been made to develop benchmarks to assess the use of “LLMs as Planners.” PlanBench Valmeekam et al. (2023a) evaluates LLMs on various planning-related tasks, including plan generation, cost-optimal planning, plan verification, and other properties. Their work focuses on evaluating LLMs on their ability to generate plans. In contrast Planetarium focuses on evaluating “Planner-Augmented LLMs”. Here, the goal is to determine if an LLM can successfully translate natural language descriptions of planning problems into the correct PDDL representations of those problems.
3 Preliminaries
To present Planetarium, we first introduce planning, PDDL, and scene graphs.
3.1 Classical Planning Problems
In a classical planning problem, states are fully observable, there is a finite set of actions, and transitions are deterministic. The goal is to transition from the initial state to the goal state following a transition function. We use the set-theoretic definition of classical planning problems Ghallab et al. (2004).
Definition 1.
A planning problem $P$ is denoted by the tuple $(L,S,A,\gamma,s_{i},g)$, where:
  • $L$ is a finite set of proposition symbols representing different facts about the world.
  • $S\subseteq 2^{L}$ is a set of states. Each state $s\subseteq L$ is the set of true propositions in that state.
  • $A$ represents the set of actions, where each action $a$ consists of a set of preconditions that must hold for it to be applicable and a set of effects that modify the propositions that are true after its execution.
  • The transition function $\gamma:S\times A\to 2^{S}$ models how the world changes by an action.
  • $s_{i}$ is the initial state of the world from which the problem begins.
  • $g\subseteq L$ represents the goal propositions, indicating which propositions must be true for a state to be considered a goal state. The set of goal states is $S_{g}=\{s\in S\;|\;g\subseteq s\}$.
A plan $\pi=<a_{1},a_{n},\ldots,a_{n}>$ is a sequence of actions that leads from $s_{i}$ to any state in $S_{g}$ following $\gamma$. This plan is the solution to a planning problem $P$.
3.2 Planning Domain Definition Language
The Planning Domain Definition Language (PDDL) is a specialized language designed to provide a unified way to represent planning problems. It can represent various types of planning problems, including classical planning problems McDermott (2000); Gerevini and Long (2006). A PDDL planning problem consists of two files: the domain file, which describes the constant parts of the world model, and a problem file, which specifies particular planning scenarios using the world model outlined in the domain file.
In the domain file, we specify the set of possible actions $A$ and their preconditions and effects, which collectively define the transition function $\gamma$. The problem file defines the initial state $s_{i}$ and the goal propositions $g$. Finally, the set of proposition symbols $L$ is created by combining the predicates from the domain file with the objects defined in the problem file.
Planetarium focuses on classical planning problems within the STRIPS subset of PDDL McDermott et al. (1998); Fikes and Nilsson (1971). STRIPS provides the basic grammar for describing actions by specifying a set of preconditions that must be met for an action to be applicable and a set of effects that modify the propositions that are true after the action’s execution. The expressiveness of problems defined in STRIPS is equivalent to those characterized by the set-theoretic definition (Definition 1) of classical planning problems Ghallab et al. (2004).
3.3 Scene Graphs
To compare different planning problems, we use scene graphs. A scene graph is a data structure commonly used in fields such as computer vision and graphics Johnson et al. (2015); Chang et al. (2023), rearrangement Ramachandruni et al. (2023), and planning to represent objects, their attributes, and the relationships among them. In our work, we define scene graphs as directed graphs with types and attributes for both nodes and edges. We represent a PDDL problem file with scene graphs as follows. We create one scene graph (the initial scene) for the initial state and another for the set of goal propositions (the goal scene). For every object, we create a node with an object type. Then for every proposition that is listed in the problem file, we create a node with a proposition type. That node is given an attribute with the name of its predicate. Then, for each argument to the predicate in that proposition, we add an edge from the proposition node to the corresponding object. Each such edge is given three attributes: the name of the predicate, the position of the argument (first, second, etc.), and whether it is defined in the initial state or the goal propositions. A scene graph is thus ($(\{O\cup P\},E)$, where $O$ is the set of nodes with object type, $P$ is the set of nodes with Proposition type, and $E$ is the set of edges.
We further define a problem graph as the combination of an initial scene and goal scene. Given $SceneGraph_{init}=(\{O\cup P_{init}\},E_{init})$ and $SceneGraph_{goal}=(\{O\cup P_{goal}\},E_{goal})$, then a problem graph merges the scene graphs such that $ProblemGraph=(\{O\cup P_{init}\cup P_{goal}\},\{E_{init}\cup E_{goal}\})$. We define graph isomorphism on any of these graphs as an edge and type-preserving bijection between nodes, meaning that two graphs share a connectivity structure where all types and attributes match. See Appendix B for diagrams of scene and problem graphs.
4 Evaluation Framework
In this section, we describe the design of Planetarium. Our goal is to systematically evaluate how well LLMs can translate natural language descriptions of planning tasks into structured planning languages. Planetarium consists of two components: an algorithm that validates whether a ground truth PDDL problem file and an LLM-generated PDDL problem file are equivalent, and a curated dataset of planning tasks against which an LLM can be evaluated.
4.1 Planning Problem Equivalence
The first step to benchmarking PDDL generation is determining how to decide whether the generated code matches ground truth code. One might assume that checking if two PDDL problem files are equivalent is straightforward. However, the same initial state, goal state pair could be represented by many PDDL problem files, as shown in Figure 1.
Given these difficulties, we propose a definition of equivalence in terms of classical planning problems. The main idea is to find a bijective function between the sets of proposition symbols $L$ of each problem, such that when applied to the other elements, this function makes both problems equal. This definition assumes that the set of actions $A$ and the transition function $\gamma$ are shared between the two problems. Our formal definition of equivalence between two planning problems is:
Definition 2.
Two planning problems $P_{1}=(L^{1},S^{1},A,\gamma,s_{i}^{1},g^{1})$ and $P_{2}=(L^{2},S^{2},A,\gamma,s_{i}^{2},g^{2})$ with the same actions $A$ and transition function $\gamma$ are equivalent if there exists a bijective function $f:L^{1}\to L^{2}$ such that:
  1. 1.
    $S^{2}=\{\{f(p):p\in s\}:s\in S^{1}\}$
  2. 2.
    $s_{i}^{2}=\{f(p):p\in s_{i}^{1}\}$
  3. 3.
    $S_{g^{2}}=\{\{f(p):p\in s\}:s\in S_{g^{1}}\}$
This definition is not directly usable for checking equivalence between two PDDL problem files. This is because our definition relies on finding a bijection between the sets $L$ of each problem, and PDDL does not define these sets directly. To build the set $L$ with PDDL, one needs to use the predicates in the domain file and the objects in the problem file, instantiating each predicate with all possible objects they might take. Since the set of actions $A$ and the transition function $\gamma$ are defined in the PDDL domain file, and since the equivalence definition assumes they are shared, we can assume the entire PDDL domain is shared. This fact entails that the predicates will be shared, making it necessary to look for bijective functions for PDDL only over the objects. The challenge is that each PDDL problem file can correspond to many pairs of initial and goal states because the set of goal propositions makes an open-world assumption. That set can leave implicit trivial propositions that are necessarily true and therefore do not change the underlying planning problem. We must therefore fully specify the goal, meaning that we identify all propositions that are true in all reachable goal states when starting from the initial state.
Our algorithm for checking equivalence is summarized in Algorithm 1. First, we transform each set of initial state and goal state propositions into scene graphs. Second, we fully specify the goal scene graphs by adding all trivially true edges. Finally, we join the initial state scene graph with each goal state graph to create problem graphs and look for a bijection between objects such that the problem graphs are isomorphic. We now discuss these components in more detail.
Algorithm 1 Planning Problem Equivalence
1:function isEquivalent($\text{P}_{a},\text{P}_{b},\text{isPlaceholder}$)
2:     $\text{s}_{i,a},\text{g}_{a}\leftarrow\text{toSceneGraph}(\text{P}_{a})$
3:     $\text{s}_{i,a},\text{g}_{a}\leftarrow\text{toSceneGraph}(\text{P}_{b})$
4:
[…]
6:         return $\text{fastEquivalence}(\text{s}_{i,b},\text{g}_{a},\text{s}_{i,a},\text{g}_{b})$
7:     end if
8:
9:     $\text{g}^{\star}_{a}\leftarrow\text{fullySpecify}(\text{s}_{i,a},\text{g}_{a})$
10:     $\text{g}^{\star}_{b}\leftarrow\text{fullySpecify}(\text{s}_{i,b},\text{g}_{b})$
11:
12:     if isPlaceholder then
13:         return $\text{isIsomorphic}(\text{s}_{i,a},\text{s}_{i,b})\wedge\text{isIsomorphic}(\text{g}^{\star}_{a},\text{g}^{\star}_{b})$
14:     else
15:         $\text{p}_{a}\leftarrow\text{join(}\text{s}_{i,a},\text{g}^{\star}_{a}\text{)}$
16:         $\text{p}_{b}\leftarrow\text{join(}\text{s}_{i,b},\text{g}^{\star}_{b}\text{)}$
17:
18:         return $\text{isIsomorphic}(\text{p}_{a},\text{p}_{b})$
19:     end if
20:end function
Transform to Scene Graphs. From each PDDL problem file, we generate two scene graphs: one for the initial state and another for the goal state (lines 2–3). For each transformation, we first create an object node for each object. Then, for each proposition in the collection, we create a new proposition node and create edges between the proposition node and the objects it takes as arguments. Edge attributes denote argument order, predicate type, and whether the proposition is in an initial or goal scene. See Section 3.3 for further details.
Check Easy Cases. For speed, we check several cases and return early if we can (lines 5–7). First, if the number of object nodes in each graph is not equal, the problems cannot be equivalent. Second, if the initial scenes are not isomorphic, then the problems cannot be equivalent. Finally, if the problem graphs, composed of the initial and goal scenes, are isomorphic, then the problems are equivalent.
Fully Specify the Goal Scenes. If the input is not an easy case, then we have to reason about the sets of goal states defined by the goal scenes. Condition 3 of Definition 2 requires that the sets of goal states that are consistent with the given goal propositions be equal after substituting matching propositions. Since PDDL uses the open-world assumption for goals, we have to identify all propositions that are true in all reachable goal states when starting from the initial state for each problem. The function fullySpecify finds all such propositions and adds them to the goal scenes as additional proposition nodes with the corresponding edges and attributes (lines 9–10). We show how to implement fullySpecify efficiently for Blocks World and Gripper in Appendix C.
PDDL without Object Identity. Our algorithm runs in two modes, depending on the type of problem we are checking. Sometimes we want to compare a generated PDDL problem file with many ground truth problem files and see if the generated file matches any one of them. For example, if the natural language description says to “make a tower of height 3,” the specific blocks to use are unspecified, and any permutation of blocks that builds a tower of height 3 should be considered correct. Concretely, we want to treat the objects in the PDDL goal states as placeholders and accept any permutation of them. We check this condition when isPlaceholder is True (determined by the problem type) simply by checking isomorphism between initial and goal scenes separately rather than combining them into problem graphs (line 13).
PDDL with Object Identity. If isPlaceholder is False then we want to check whether the problem files are precisely equivalent under Defintion 2, meaning that the objects in the initial scenes correspond to the same objects in the goal scenes as well. We first join the corresponding initial and scene graphs into problem graphs as described in Section 3.3 (lines 15–16). Then we check whether the two problem graphs are isomorphic (line 18).
We illustrate Algorithm 1 with examples in Appendix B. We also formally state its correctness, with the proof in Appendix A.
Theorem 1.
isEquivalent($P_{a}$, $P_{b}$, False) returns True if and only if the PDDL problem files $P_{a}$ and $P_{b}$ represent equivalent planning problems under Definition 2. isEquivalent($P_{a}$, $P_{b}$, True) returns True if and only if $P_{a}$ represents a planning problem that is equivalent to some planning problem represented by $P_{b}$ after a permutation of the objects in its goal state.
4.2 The Dataset
The Planetarium dataset includes 132,037 text-to-PDDL pairs across 13 tasks, derived from the Blocks World and Gripper domains Seipp et al. (2022); McDermott (2000).
Dataset Construction.
We focus on these two domains as they are commonly used in other works Valmeekam et al. (2023a); Liu et al. (2023), but are nevertheless challenging. Each text-to-PDDL pair consists of a prompt detailing the initial and goal states, along with the corresponding ground truth PDDL. (See Appendix D for examples.) Table 1 shows a breakdown of the Planetarium data by text description abstractness and number of propositions. Each instance is a ground truth PDDL problem description and text description pair.
Table 1: A breakdown of the Planetarium dataset by the level of abstractness of the text description and number of propositions in the ground truth PDDL problem (size). Each instance is a ground truth PDDL problem description and text description pair.
Domain Total Abstractness Num. of Ground Truth Propositions (Size)
Explicit to Abstract to
Explicit Abstract Explicit Abstract 1-20 21-40 41-60 61-80 >80
Blocks World 92,350 23,033 23,087 23,086 23,144 1,012 10,765 50,793 26,316 3,464
Gripper 39,687 10,632 9,518 10,313 9,224 379 2,112 9,412 25,346 2,438
We vary the data along two dimensions: abstractness (explicit vs abstract) and size. Explicit planning problem text descriptions correspond directly to propositions found in the problem PDDL (e.g., “block 1 is on block 2”). Abstract text descriptions instead summarize a state (e.g., “all blocks are in a single tower”). Since our text descriptions describe both initial and goal states, each can be either an abstract or explicit description, leading to four possible categories: explicit to explicit, explicit to abstract, abstract to explicit, and abstract to abstract. We measure the size of a problem by the number of propositions listed in the ground truth problem PDDL. Larger problems typically pose greater challenges for LLMs.
To obtain interesting problems for each domain, we crafted a set of tasks instead of randomly generating instances. Below are descriptions of each domain and their corresponding tasks.
Blocks World. This domain involves using a robotic hand to manipulate a set of blocks arranged in various configurations on a table. We have designed the following tasks for this domain:
  • Stack: Arrange all the blocks into a single tower.
  • Unstack: Dismantle any tower so that all blocks are separately placed on the table.
  • Holding a Block: Similar to unstacking, but one block remains held by the robot.
  • Staircase: Create stacks of increasing sizes from one upwards until no blocks are left.
  • Equal Towers: Arrange the blocks into towers of equal height.
  • Swap: With two towers, swap the base blocks and leave the rest of the structure unchanged.
  • Invert: Invert each stack so the entire sequence of blocks is reversed.
  • Towers: Build towers to match a specified list of heights.
Gripper. This domain features a robot equipped with grippers, tasked with transporting balls from one room to another. We have designed the following tasks for this domain:
  • Single Room: Move all balls to one room.
  • Evenly Distribute: Position the balls so each room has the same number of balls.
  • Arbitrarily Distribute: Move the balls according to a predefined list that specifies the number of balls per room.
  • Swap Rooms: Exchange the balls between two rooms.
  • Move to Max: Move all balls to the room currently containing the most balls.
  • Move to Min: Move all balls to the room with the least number of balls.
5 Evaluating LLMs on Planetarium
Figure 2: Performance of various models on the Planetarium test set.
As an initial snapshot of the field’s current state, we evaluate several API-access and open-weight language models on Planetarium in both zero-shot and fine-tuned settings. We find that while powerful models like GPT-4o can often generate valid PDDL problems (in the sense that some plan solves them), they are rarely correct. This result underscores the need for Planetarium’s rigorous approach to evaluation of PDDL generation. The code to recreate the entire evaluation is available at https://github.com/BatsResearch/planetarium.
Models. We evaluate one API-access model (GPT-4o) and three open-weight models (Mistral v0.3 7B Instruct Jiang et al. (2023), and Gemma 1.1 IT 2B & 7B Team et al. (2024)).
Evaluation Protocol.
We evaluate models on the Planetarium test set, which we built by randomly splitting the Planetarium dataset with an $80/20$ train/test split. Models are prompted with the natural language description of the task along with the respective domain PDDL. We measure three metrics: the number of syntactically parseable problems generated, the number of solve-able problems, and the number of semantically correct problems. We say a model output is parseable if a PDDL parser supporting :strips can parse a valid PDDL problem from a substring in the output and if it can be converted into our graph representation. A problem is solve-able if it is parseable and there exists a plan that can be applied to the initial scene that results in the goal scene. Due to the size and complexity of some of the problems in our dataset, a generalized classical planner cannot always reliably and quickly return solutions. Therefore, we write specialized planners that work for all problems in our dataset, and generally, all validly defined blocksworld and gripper domain problems, except a few invalid edge cases (e.g., one block on top of two blocks at a time, holding two blocks, etc.). To evaluate the correctness of model outputs, they must be first parseable and solve-able. Then, we use our PDDL equivalence algorithm to verify equivalence to our ground truth PDDL. We found our equivalence algorithm to take on average $12$ms per example to compute on an M2 Apple Silicon laptop with batch parallelization.
Fine-Tuning. We fine-tuned the open-weight models using QLoRA with a rank of 16, adhering to the hyperparameter recommendations for small models provided by the original authors Dettmers et al. (2023). Models were loaded in 4-bit precision and were trained over the training set for a single epoch. Fine-tuning for each model used either two NVIDIA GeForce RTX 3090 GPUs or two NVIDIA A6000 GPUs, operating with data parallelization for approximately 15 hours. We truncate the longest $5\%$ of our training dataset due to GPU memory constraints. Experiments were conducted using a GPU cluster at the Center for Computation and Visualization, Brown University. Additional experiment details are provided in Appendix F.
Results.
The performance of GPT-4o, Mistral v0.3 7B Instruct, and Gemma 1.1 IT 2B & 7B on the Planetarium test set is shown in Figure 2. We report the percentage of generated plans that were parseable and correct and report results from zero-shot and fine-tuned settings.
Figure 3: Breakdown of GPT-4o zero-shot performance at varying levels of abstraction.
Analysis and Discussion. All models exhibited poor zero-shot performance, with GPT-4o achieving the highest accuracy. A breakdown of GPT-4o’s zero-shot performance, the only model to achieve significant zero-shot performance, shows abstract task descriptions are harder to translate than explicit ones, and fully explicit task descriptions make generating parseable PDDL code easier (Figure 3).
Fine-tuning improved performance across all open-weight models, with Mistral v0.3 7B Instruct achieving the highest accuracy. More experiments and analyses are shown in Appendix G.
6 Conclusion, Limitations, and Future Work
Planetarium is a new benchmark for assessing the ability of LLMs to translate natural language descriptions of planning problems into PDDL. A potential societal impact of this research is ensuring the correctness of translating natural language into structured planning languages. If this translation method becomes widespread but its evaluation remains inaccurate, systems could produce misleading or misaligned results that could cause harm if acted upon. Planetarium highlights the importance of assessing the correctness of translations. We achieve this by reasoning about PDDL semantics and the inherent structure of classical planning problems.
Planetarium has a few important limitations: Planetarium currently only supports the Blocks World and Gripper domains. While these domains have been popular for studying LLMs and their relation to planning, incorporating more expressive domains in the future will widen the scope of Planetarium. This benchmark is also currently restricted to the STRIPS subset of PDDL. Extending it to support more expressive subsets of PDDL will allow us to evaluate more complex, real-world planning problems such as non-deterministic, temporal, and numeric domains. Our work reveals generating semantically correct structured planning language descriptions to be a challenging task for language models, with models like GPT-4o achieving $35.1\%$ zero-shot accuracy but producing valid, seemingly correct descriptions $82.2\%$ of the time. We hope that Planetarium will drive progress on hybrid approaches combining LLMs and classic planners, setting a standard for the evaluation of such tasks.
Acknowledgments and Disclosure of Funding
This research is supported in part by the Office of Naval Research (ONR) award N00014-20-1-2115. We gratefully acknowledge support from Cisco, Cognex, and the Brown Computer Science Faculty Innovators Fund. Disclosure: Stephen Bach is an advisor to Snorkel AI, a company that provides software and services for data-centric artificial intelligence.
Appendix A Proof of Theorem 1
Proposition 1.
isEquivalent($P_{a}$, $P_{b}$, False) returns True if and only if the PDDL problem files $P_{a}$ and $P_{b}$ represent equivalent planning problems under Definition 2.
Proof.
To prove this proposition, we need to prove each direction of the biconditional. First, we will prove the forward implication. There are two possible ways the algorithm can return True: Either in the isIsomorphic check (line 18) or in fastEquivalence (line 6).
  1. Case 1:
    When isIsomorphic($\text{p}_{a}$, $\text{p}_{b}$) returns True, we know that there exists a bijection $\varphi$ from vertices $V_{p_{a}}$ to vertices $V_{p_{b}}$. Since $\varphi$ is a type-preserving bijection, it means the two graphs share a connectivity structure where all types and attributes match.
    By using Lemma 1, we can build a new bijection $f$ from $\varphi$. Instead of mapping vertices in the scene graph, $f$ maps propositions from $L^{a}$ in planning problem $P_{a}$ to propositions in $L^{b}$ in planning problem $P_{b}$. Given that $\varphi$ is a type-preserving bijection, we know that the bijection $f$ derived from $\varphi$ also maps objects to objects, predicates to predicates, vertices in the initial scene $s_{i,a}$ to vertices in $s_{i,b}$, and vertices in the goal scene $g_{a}^{*}$ to vertices in $g_{b}^{*}$.
    To prove Property 1, we can build the set $L^{a}$ for each planning problem by taking each predicate in the PDDL domain and instantiating it with all possible combinations of objects from the problem PDDL $P_{a}$. Then, we can construct the set of all possible states of the problem by taking the power set $S^{a}=2^{L^{a}}$. For each proposition $p\in s$ in each state $s\in S^{a}$, we apply the bijection $f$ to show that we get $S^{b}$.
    $S$ =$=\{\{f(p):p\in s\}:s\in S^{a}\}$ (1)
    $S$ =$=\{\{\langle\varphi(q^{a}),\varphi(o_{1}^{a}),\varphi(o_{2}^{a}),\ldots,\varphi(o_{k}^{a})\rangle:p\in s\}:s\in S^{a}\}$ (2)
    $S$ =$=\{\{\langle q^{b},o_{1}^{b},o_{2}^{b},\ldots o_{k}^{b}\rangle:p\in s\}:s\in S^{a}\}$ (3)
    $S$ =$=\{\{p_{b}:p\in s\}:s\in S^{a}\}$ (4)
    $S$ $=S^{b}$ (5)
    To prove Property 2, we go through each proposition $p$ in the initial state $s_{i}^{a}$ of the problem PDDL $P_{a}$, and apply the bijection $f$ to show that we get $s_{i}^{b}$.
    $s_{i}$ =$=\{f(p_{a}):p_{a}\in s_{i}^{a}\}$ (6)
    $s_{i}$ =$=\{\langle\varphi(q^{a}),\varphi(o_{1}^{a}),\varphi(o_{2}^{a}),\ldots,\varphi(o_{k}^{a})\rangle:p_{a}\in s_{i}^{a}\}$ (7)
    $s_{i}$ =$=\{\langle q^{b},o_{1}^{b},o_{2}^{b},\ldots o_{k}^{b}\rangle:p_{a}\in s_{i}^{a}\}$ (8)
    $s_{i}$ =$=\{p_{b}:p_{a}\in s_{i}^{a}\}$ (9)
    $s_{i}$ $=s_{i}^{b}$ (10)
    To prove Property 3, we first observe that the fullySpecify function adds all possible trivial propositions to the goal scene graph. By definition, trivial propositions are those that are known not to change the underlying planning problem when added or removed. We can use the same bijection $f$ constructed from $\varphi$. We don’t need to add or remove non-trivial predicates, as doing so would change the planning problem by definition. This implies that the augmented goal state represents any other possible goal state in $S_{g}$. With this, we can use the bijection $f$ built from $\varphi$ to apply it to any state in $S_{g^{a}}$ and show that we get $S_{g^{b}}$.
    $S_{g}$ =$=\{\{f(p):p\in s\}:s\in S_{g^{a}}\}$ (11)
    $S_{g}$ =$=\{\{\langle\varphi(q^{a}),\varphi(o_{1}^{a}),\varphi(o_{2}^{a}),\ldots,\varphi(o_{k}^{a})\rangle:p\in s\}:s\in S_{g^{a}}\}$ (12)
    $S_{g}$ =$=\{\langle q^{b},o_{1}^{b},o_{2}^{b},\ldots o_{k}^{b}\rangle:p\in s\}:s\in S_{g^{a}}\}$ (13)
    $S_{g}$ =$=\{\{p_{b}:p\in s\}:s\in S_{g^{a}}\}$ (14)
    $S_{g}$ $=S_{g^{b}}$ (15)
  2. Case 2:
[…]
    In Case 1 we have proven that when isIsomorphic($p_{a}$, $p_{b}$) returns True, the planning problems $P_{a}$ and $P_{b}$ are equivalent under the bijection $f$. Since $p_{a}=\texttt{join(}s_{i,a},g_{a}^{\star}\texttt{)}$ and $p_{b}=\texttt{join(}s_{i,b},g_{b}^{\star}\texttt{)}$, we can let $g_{a}^{\star}=g_{a}$ and $g_{b}^{\star}=g_{b}$. This is because, by definition, trivial propositions do not change the underlying planning problem when added or removed, and in this case, we are only removing trivial propositions. Hence, by the proof of Case Case 1: we conclude that this case has also to be true.
Now, we will prove the backward implication. In this case, we will assume that there exists a bijection $f$ that follows properties 12, and 3 of Definition 2. By Lemma 1, we know that $f$ comes from $\varphi$, and since both are bijections, we can recover $\varphi$ by doing the inverse process. We need to prove that these conditions lead to isIsomorphic (line 18) or fastEquivalence (line 6) returning True.
Since there exists a bijection $\varphi$ that comes from $f$, the number of objects must be the same for planning problems $P_{a}$ and $P_{b}$, as stated in Property 1. With Property 2, we know that this bijection $\varphi$ must also make the initial scenes isomorphic. Assume that the sets of goal propositions have the same size. By using Property 3, we know that the bijection $\varphi$ would make fastEquivalence (line 6) return True.
Conversely, assume that the sets of goal propositions have different sizes. In this case, canFastEquivalence would not allow fastEquivalence to execute. In the case of isIsomorphic (line 18), we know that $p_{a}=\texttt{join(}s_{i,a}g_{a}^{\star}\texttt{)}$ and $p_{b}=\texttt{join(}s_{i,b},g_{b}^{\star}\texttt{)}$. Since we know that we can construct the bijection $\varphi$ that maps vertices from the bijection $f$, we also know that since $\varphi$ comes from $f$, this bijection respects initial states (Property 2) and goal states (Property 3). Therefore, isIsomorphic($p_{a}$, $p_{b}$) must return True.
With this, we have proven that if two problems are equivalent under Definition 2, then the equivalence algorithm must return True.
Since we have proven both directions of the biconditional, Proposition 1 is true. ∎
Proposition 2.
isEquivalent($P_{a}$, $P_{b}$, True) returns True if and only if $P_{a}$ represents a planning problem that is equivalent (under Definition 2) to some planning problem represented by $P_{b}$ after a permutation of the objects in its goal state.
Proof.
To prove this proposition, we need to prove each direction of the biconditional. First, we will prove the forward implication. There are two possible ways the algorithm can return True: Either in fastEquivalence (line 6) or in isIsomorphic($s_{i,a}$, $s_{i,b}$) $\wedge$ isIsomorphic($g^{\star}_{a}$, $g^{\star}_{b}$).
To prove the case where fastEquivalence (line 6) returns True, we can use the same argument as in Case Case 1: of the forward implication in the proof of Proposition 1. This is because the placeholder being True does not affect the fastEquivalence function.
Now, assume that isIsomorphic($s_{i,a}$, $s_{i,b}$) $\wedge$ isIsomorphic($g^{\star}_{a}$, $g^{\star}_{b}$) is True. This implies that there exist two bijections: one $\varphi$ on the vertices $V_{s_{i,a}}$ to $V_{s_{i,b}}$, and another $\psi$ between vertices $V_{g^{\star}_{a}}$ to $V_{g^{\star}_{b}}$. Now, by using Lemma 1, we can build two new bijections, $f$ and $h$, that map between propositions in $L$ instead of nodes in the scene graph, as $\varphi$ and $\psi$ do.
Since the underlying mappings of $f$ and $h$ can be different, we need a way to combine them into a single bijection. One way to construct this bijection is to permute the mappings of $f$ until they fit the mappings of $h$. This implies that there are many pairs of $f$ and $h$, and each pair can be combined into a global bijection using function composition $f\circ h$. For each of these permutations, we can then use the arguments in Case Case 2: of the proof of Proposition 1 to show that these pairs hold all three properties of Definition 2.
Now, we will prove the backward implication. First, since fastEquivalence remains the same, we do not need to prove anything new about it; we can simply reuse the previous results.
Assume $P_{a}$ represents a planning problem that is equivalent to some planning problem represented by $P_{b}$ after a permutation of the objects in its goal state. This implies that there are several bijections $f$ from the predicates in $P_{a}$ to the predicates in $P_{b}$. For each of these bijections $f$, we can use Lemma 1 to transform it into $\varphi$. Since we know by Property2 that $\varphi$ maps initial states, there is an isomorphism between the initial states. By Property 3, we know that $\varphi$ maps any goal state of $P_{a}$ to any other possible state in $P_{b}$. Hence, $s^{\star}_{a}$ has to be isomorphic with $s^{\star}_{b}$ since $\varphi$ is a bijection between their nodes.
Since we have proven both directions of Proposition 2, we know it is true. ∎
Proof of Theorem 1.
Having the proof of Proposition 1 and the proof of Proposition 2, we know that the theorem, which is just the conjunction of these propositions, must be true. ∎
Lemma 1.
Let $\varphi$ be an isomorphic bijection between nodes in two scene graphs (propositions or objects). There exists a bijection $f$ over the propositions in $L$ such that for any proposition $p\in L$, $f(p)=\langle\varphi(q),\varphi(o_{1}),\varphi(o_{2}),\ldots,\varphi(o_{k})\rangle$, where $q$ is the predicate and $O=\{o_{1},o_{2},\ldots,o_{k}\}$ is a set of objects.
Proof.
Take any proposition $p\in L$ and identify its predicate $q$ and the ordered set of objects it acts on, $O=\langle o_{1},o_{2},\ldots,o_{k}\rangle$. Then, we can transform the proposition $p$ into the ordered tuple $p=\langle q,o_{1},o_{2},\ldots,o_{k}\rangle$. Finally, we can build $f(p)=\langle\varphi(q),\varphi(o_{1}),\varphi(o_{2}),\ldots,\varphi(o_{k})\rangle$, which is a bijection over the propositions in $L$. ∎
Appendix B Examples of Algorithm 1
Figure 4: An illustration of the algorithm to check if two PDDL problems are equivalent. It shows each of the stages of the algorithm: transforming to scene graphs, fully specifying the goal propositions, and checking for graph isomorphism.
Appendix C Implementation of fullySpecify
The fullySpecify function adds trivial edges (edges that must exist but are not currently present) to our scene graphs. We build the fullySpecify function using our domain knowledge of the Blocks World and Gripper domains.
Blocks World
The following helpful facts are true about Blocks World:
  • If all blocks have its behavior above it defined (they are either clear, have something on top of it, or is being held), then any block that does not have its bottom behavior specified, must be on the table (on-table).
  • If all blocks have its behavior beneath it defined (they are either on top of something else, are on the table (on-table), or are being held), then any block that does not have its behavior above it defined must be clear.
  • Following the last two facts: if there is a block A which has its behavior above it unspecified and block B which has its behavior below it unspecified, then the only way that (on B A) can exist is if there isn’t already a chain of predicates that leads A to B or vice versa. If there already exists a chain connecting these two blocks, that means A and B are already connected by being in the same tower, and that they are the top and bottom of the tower they are a part of.
  • If there are no “floating” blocks (blocks that have both its top and bottom behavior defined), and the arm’s behavior is undefined (neither arm-empty nor holding) is present, then the arm must be empty. This is because there is no block that the arm could possibly hold here.
Using these rules, we can add all possible trivial edges, until no more trivial edges can be found. Further, these facts will discover all possible trivial edges in Blocks World, meaning the fully specified scene graph will be in its canonical form after fullySpecify.
Gripper
Similarly, we can build a set of facts that operate on the Gripper domain:
  • If all balls are assigned a room (at ball room), then all unassigned grippers (no (free gripper) or (carry ball gripper)) must be free.
  • If there is only one room, and if all grippers are already specified (either (free) or (carry ball gripper)), then any unspecified balls must be in the only room that exists.
  • If there is only one room, the robby must be in it (at-robby onlyRoom).
These are the only rules we can find. This is because, if a ball’s position is unspecified, it can always be assigned to any arbitrary room. If the robby is unspecified (no at-robby), then it too, can always be specified to any arbitrary room.
Appendix D Text-to-PDDL Pair Example
BLOCKSWORLD Ground Truth PDDL (define (problem equal_towers_to_equal_towers_5) (:domain blocksworld) (:requirements :strips) (:objects b1 b2 b3 b4 b5) (:init (arm-empty) (clear b5) (on b2 b1) (on b3 b2) (on b4 b3) (on b5 b4) (on-table b1)) (:goal (and (arm-empty) (on-table b1) (on b2 b1) (on b3 b2) (on b4 b3) (on b5 b4) (clear b5))) ) Natural Language Description (objects: You have 5 blocks, b1 through b5, stacked into 1 towers of equal heights, and your arm is empty. ) (abstract init: stacked into 1 towers of equal heights, and your arm is empty. | explicit init: Your arm is empty. b1 is on the table. b2 is on b1. b3 is on b2. b4 is on b3. b5 is on b4. b5 is clear. ) Your goal is to have the following: (abstract goal: stack the blocks into 1 towers of equal heights. | explicit goal: Your arm should be empty. b1 should be on the table. b2 should be on b1. b3 should be on b2. b4 should be on b3. b5 should be on b4. b5 should be clear. )
Appendix E License Information
The data for Planetarium is available at https://huggingface.co/datasets/BatsResearch/planetarium and released under a Creative Commons CC-BY-4.0 license. In addition, all code to create the dataset and evaluate models on the benchmark is available at https://github.com/BatsResearch/planetarium and released under a BSD 3 license. The authors are responsible for the content of the dataset.
In our study, we use the Blocksworld and Gripper domains, both of which are sourced from the IPC McDermott [2000]. The specific domain files used in our research are obtained from pddl-generators  Seipp et al. [2022], which is distributed under the GNU General Public License.
Appendix F Additional Experiment Details
Table 2 displays the the hyperparameters used for fine-tuning across all models. In addition, models are loaded using 4-bit NF4 quantization with double quantization to reduce the average memory footprint. We also use bfloat16 compute data type for faster training.
Hyperparameters Value
Optimizer adamw_torch
Learning rate 2e-5
Batch Size 1
Betas (0.9, 0.999)
Epsilon 1e-8
Weight Decay 0.01
Max Sequence Length 1500
LoRA rank 16
LoRA alpha 32
LoRA Dropout 0.05
Table 2: Training hyperparameters when fine-tuning all models included in the paper. The hyperparameters are seperated in three parts: supervised fine-tuning related, generation related, and LoRA related.
Appendix G Additional Experiments and Analysis
Figure 5: (Left) Gemma 1.1 IT 2B, fine-tuned on the random train split and evaluated on certain tasks. Most tasks are performed extremely well. (evaluated tasks were seen during training). (Right) Gemma 1.1 IT 2B, fine-tuned on all examples except for the tasks listed above, along with one_room in the initial state and n_room_distributed in the initial state. (tasks not seen during training)
Appendix H Dataset Documentation
To see the most up-to-date documentation, go to the following URL:
https://datanutrition.org/labels/v3/?id=f2b30848-2b93-4f2d-b7e0-926604a02c12
Appendix I Maintenance Plan
The BATS research group at Brown University, will maintain the Planetarium dataset and benchmark with an open GitHub repository and issue submission system, and the dataset hosted on HuggingFace. The maintenance plan includes regular issue tracking, with reviews and categorization of issues, aiming to resolve high-priority issues within a week and scheduling minor updates quarterly. Comprehensive documentation and automated testing will ensure quality and compatibility. Francisco Piedrahita Velez and Max Zuo will lead the maintenance efforts, supported by the BATS research group.